OpenStudy (anonymous):

Partial Fraction Decomposition...x^(2)+1/x^(3)+x^(2)-5x+3

OpenStudy (anonymous):

\[\frac{x^2+1}{x^3+x^2-5x+3}\]?

OpenStudy (anonymous):

this is going to suck because \[x^3+x^2-5x+3=(x-1)^2(x+3)\]

OpenStudy (anonymous):

you have to write \[\frac{x^2+1}{x^3+x^2-5x+3}=\frac{A}{x-1}+\frac{B}{(x-1)^2}+\frac{C}{x+3}\]and then find the coefficients up top

OpenStudy (anonymous):

Okay! I was just wondering where to put x+3 I was putting it with the A previously. Thanks!!

OpenStudy (anonymous):

it makes no difference what you label the numerators that is not the problem the problem is finding them

OpenStudy (anonymous):

Oh okay! I hope I can figure it out. Haha!

OpenStudy (anonymous):

i can help with that if you like

OpenStudy (anonymous):

Sure, if you don't mind

OpenStudy (anonymous):

when you add that mess on the left, the numerator will be \[A(x-1)(x+3)+B(x+3)+C(x-1)^2\] so first set \[x^2+1=A(x-1)(x+3)+B(x+3)+C(x-1)^2\]

OpenStudy (anonymous):

i meant on the right doh

OpenStudy (anonymous):

we can find \(B\) instantly by replacing \(x\) by \(1\) and getting \[2=4B\] so \[b=\frac{1}{2}\]

OpenStudy (anonymous):

How come there isn't a (x-1) with the B??

OpenStudy (anonymous):

because the denominator of B is \((x-1)^2\)already, so all you need to do to that term to put it over the common denominator \((x-1)^2(x+3)\) is multiply top and bottom by \(x+3\)

OpenStudy (anonymous):

Okay that makes sense. It gets cancelled out basically?

OpenStudy (anonymous):

? nothing is cancelled

OpenStudy (anonymous):

So you don't even multiply it with x-1 ?

OpenStudy (anonymous):

the expression i wrote on the right of \[x^2+1=A(x-1)(x+3)+B(x+3)+C(x-1)^2\] is the numerator if you were to add \[\frac{A}{x-1}+\frac{B}{(x-1)^2}+\frac{C}{x+3}\]

OpenStudy (anonymous):

to add, you would have to multiply \(\frac{A}{x-1}\) top and bottom by \((x-1)(x+3)\) because the denominator will be \((x-1)^2(x+3)\)

OpenStudy (anonymous):

you would have to multiply \(\frac{B}{(x-1)^2}\) top and bottom by \(x+3\) and you would have to multiply \(\frac{C}{x+3}\) top and bottom by \((x-1)^2\) that would put each term over the right denominator

OpenStudy (anonymous):

that is why i wrote \[x^2+1=A(x-1)(x+3)+B(x+3)+C(x-1)^2\]

OpenStudy (anonymous):

so far so good?

OpenStudy (anonymous):

yeah! thanks

OpenStudy (anonymous):

ok now did you get how i got \(B=\frac{1}{2}\)?

OpenStudy (anonymous):

Yeah!

OpenStudy (anonymous):

now comes the hard part

OpenStudy (anonymous):

oh maybe not lets try making \(x=-3\)

OpenStudy (anonymous):

then \[x^2+1=A(x-1)(x+3)+B(x+3)+C(x-1)^2\] turns in to \[10=16C\] so \[C=\frac{5}{8}\]

OpenStudy (anonymous):

that ok too?

OpenStudy (anonymous):

Yep!

OpenStudy (anonymous):

ok now comes the think part, because i really don't want to multiply all that crap out, which is another way to do it, but i don't like doing that if i can avoid it, because i often make an algebra mistake

OpenStudy (anonymous):

visualize the constant when you multiply this out \[A(x-1)(x+3)+B(x+3)+C(x-1)^2\] it should be clear it is \[-3A+3B+C\] right?

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

sure? if not, you have to multiply all that crap out

OpenStudy (anonymous):

I actually didnt get the +C?

OpenStudy (anonymous):

but lets pretend it is more or less obvious now we know \(B=\frac{1}{2}, C=\frac{5}{8}\)

OpenStudy (anonymous):

Wait nevermind..got it!

OpenStudy (anonymous):

k good it is the constant when you square \((x-1)\)

OpenStudy (anonymous):

on the left, the constant is 1 on the right, the constant is \[-3A+3B+C\] i.e. \[1=-3A+3\times \frac{1}{2}+\frac{5}{8}\] solve for A

OpenStudy (tkhunny):

"we can find B instantly by replacing x by 1 and ..." Keep in mind that this may be confusing. Fundamentally, x = 1 is NOT in the Domain of the Original Function. This odd substitution makes little sense, excepting that it does help to find the solutions more quickly in many cases.

OpenStudy (anonymous):

true, but i was not putting it in the original function actually there is an even snappier way to do it but i am not sure i can draw it here

OpenStudy (anonymous):

in any case if we did it right we should get \(A=\frac{3}{8}\) and that is the end

OpenStudy (anonymous):

I was taught this weird way of doing with a system of equations and then you take out the x^2 and x. I ended up with the answers A=-1/2 B=-2/3 C=3/2 and wow what a difference! thanks how would I write what we got?

OpenStudy (tkhunny):

It doesn't matter where you put it. It NOT in the Domain. You should not be putting it anywhere. It still works, though.

OpenStudy (anonymous):

Do you know which of the answers is correct? I am so confused!

OpenStudy (tkhunny):

You should not be confused. Just drag through it and set it up and solve it. Final Answer: \(\dfrac{3/8}{x-1}+\dfrac{1/2}{(x-1)^{2}}+\dfrac{5/8}{x+3}\)